# SWIFT: A Monotonic, Flux-Form Semi-Lagrangian Tracer Transport Scheme for Flow with Large Courant Numbers

### v1

@@ 1.1 Desirable Properties for Transport Schemes
Numerical methods for solving the transport equation are a core component of atmospheric models that are used for weather prediction and climate research. These equations typically take two forms: the conservative form (1a) for a density field $\rho$ or the advective form (1b) which may be used for instance for a tracer mixing ratio [^1] The mixing ratio of a species is defined as the ratio of the mass density of the species to some reference mass density, which in this work is denoted by $\rho$. In atmospheric applications this could be the density of dry air. $m$:
@@
where $\bm{u}$ is the wind velocity field. Considering the equation for the product $\rho m$ (which describes the density of the tracer species), it can be shown that this obeys:
@@
The relevance of this form of the transport equation is that discretisations of (1a) and (2) provide local (and hence global) mass conservation, in that each cell has a closed mass budget when taking into account the mass fluxes into neighbouring cells. As discussed by Thuburn [2008], these conservation properties are desirable for weather and particularly climate models, so that the discretisation itself does not become a net source or sink of the tracer. Locally-conservative schemes naturally provide a closed budget for the tracer species. Additionally, for the transport of a density, Lin and Rood [1996] and Melvin et al. [in review] argue that any conservative transport scheme should provide increments that are linear in $\bm{\nabla\cdot u}$.
@@
However other properties are also important for tracer transport schemes: it can be vital that the scheme is positivity-preserving so that unphysical negative concentrations are not produced. A stronger possible requirement is that of monotonicity, which prevents the introduction of new minima or maxima in the solution. Note that as discussed by Lauritzen et al. [2011], in contrast to $m$, the product $\rho m$ will not in general remain bounded by its initial extrema in divergent flows, but its evolution should be shape-preserving in the sense that no new unphysical extrema are generated. In this work we focus on the monotonicity requirement of keeping the minima and maxima of mixing ratios bounded by their initial values. It is also important for a conservative tracer transport scheme to be consistent with the advective form of the equation, so that a constant mixing ratio field is preserved, independent of the flow (see [Jöckel et al., 2001] and [Zhang et al., 2008]). Finally, many operational numerical weather prediction models use large time steps. It is therefore highly attractive to use a scheme which is stable at large Courant numbers [^2] For instance, the finest resolution global configurations discussed in Walters et al. [2019] correspond to Courant numbers of around 5 in the mid-latitude jets. whilst maintaining computational efficiency and without compromising the other properties listed here. A complete discussion of properties that are desirable for transport schemes can be found in Lauritzen et al. [2011].
@@
In addition to the desirable properties listed above, the transport scheme must work within the parameters of the atmospheric dynamical core, which may result in additional requirements for the transport scheme. Many dynamical cores (e.g. [Chen et al., 2008, Wood et al., 2014, Girard et al., 2014, Melvin et al., in review]) use hexahedral cells and make use of C-grid staggering in the horizontal and Charney-Phillips staggering [Charney and Phillips, 1953] in the vertical, primarily due to the favourable representation of gravity wave dispersion (see for instance Staniforth and Thuburn [2012] and Thuburn [2022]). With these staggerings, the density $\rho$ of dry air is represented at cell centres, the wind velocity $\bm{u}$ is described at the centres of cell faces and the potential temperature $\theta$ is co-located with the vertical component of the wind velocity (at the centre of the top/bottom faces of cells). It is common for tracers such as moisture mixing ratios to be co-located with the potential temperature (or other entropy-type variable) as this is beneficial for capturing the latent heat exchanges associated with phase changes [Bendall et al., 2023]. This work considers a transport scheme that can be used with these staggerings.
@@ 1.2 Background to Flux-Form Semi-Lagrangian Schemes
Some notable conservative transport methods which meet many of the criteria above include: the CSLAM scheme of Lauritzen et al. [2010] and Harris et al. [2011]; the SLICE scheme of Zerroukat et al. [2002, 2004] and Zerroukat and Allen [2012]; the multi-moment scheme of Tang et al. [2022]; and the substepped finite volume Method-of-Lines method described in Melvin et al. [in review]. However, the focus in this work is another class of schemes that possesses many of the desirable properties: Flux-Form Semi-Lagrangian (FFSL) schemes. These schemes originate from the flux integration methods of Van Leer [1974], Colella and Woodward [1984], Carpenter Jr et al. [1990] and were developed in multiple dimensions by Leonard et al. [1996] and Lin and Rood [1996]. Thanks to their conservation and stability properties, FFSL schemes have been successfully used for transport in geophysical fluid models [Lin, 2004, Putman and Lin, 2007, Zhou et al., 2012, Neale et al., 2013, Gillibrand and Herzfeld, 2016, Harris et al., 2021, Zhang et al., 2023, Mouallem et al., 2023]. FFSL schemes involve computing mass fluxes by integrating fields between flux points and corresponding departure points. Conservation is inherent, as increments are evaluated from the divergence of fluxes, which are shared between cells. Like standard semi-Lagrangian schemes, stability is not conditional on the Courant number which permits large time steps. Instead, the stability of the scheme depends upon the Lipschitz number [^3] The one-dimensional Lipschitz number measures the one-dimensional divergence of the velocity $\bm{u}$ seen by a discretisation with time step $\Delta t$, in other words $\Delta t|\partial u/\partial x|$ in the $x$-direction and similar for other directions. For the Flux-Form Semi-Lagrangian schemes discussed here, it can be calculated for a given facet in one-dimension as $\ell=(c-c^{\dagger})\,\mathrm{sign}(c)$, where $c$ is the Courant number at that facet and $c^{\dagger}$ is the Courant number at the next facet upwind. (when this number is above 1, the trajectories of discrete particles can cross [Smolarkiewicz and Pudykiewicz, 1992]). The multi-dimensional forms of FFSL presented by Leonard et al. [1996] and Lin and Rood [1996] are built from one-dimensional calculations through dimensional-splitting techniques. Following Leonard et al. [1996], the multi-dimensional splitting of Leonard et al. [1996] and Lin and Rood [1996] will be referred to as the COSMIC splitting throughout this work. Advantages of using multiple one-dimensional methods over pure multi-dimensional methods include the ease of applying limiters, the ease of extending to large Courant number flow, and that in general one-dimensional methods are computationally cheaper than their multi-dimensional counterparts.
@@ 2.1 Review of One Dimensional Flux-Form Semi-Lagrangian Schemes
In this section, the one-dimensional FFSL calculation is reviewed, but on a general one-, two- or three-dimensional mesh. At first, the transport of a general scalar field $q$ is considered, before the scheme is applied to a density $\rho$ and a mixing ratio $m$. For simplicity, the one-dimensional scheme is described in the $x$ direction only, where $x_{i}$ are the coordinates of cell centres (with fluxes at cell facets), $i$ is the grid index, $V_{i}$ is the volume of the $i$-th cell, and $q_{i}=q(x_{i})$ is the scalar field in cell $i$. The cell facets $\varGamma_{i+1/2}$ are vertices for one-dimensional meshes, edges for two-dimensional meshes, and faces for three-dimensional meshes, and are indexed by an additional $1/2$ to indicate that these are staggered from cell centres. While $\bm{u}$ is the general wind velocity vector, $u_{i+1/2}$ is its projection onto the normal $\bm{n}$ to the facet $\varGamma_{i+1/2}$. The facet has length or area $S_{i+1/2}$, so that
$u_{i+1/2}S_{i+1/2}=u_{i+1/2}\int_{\varGamma_{i+1/2}}\bm{n}\bm{\cdot}\mathrm{d}\bm{S}=\int_{\varGamma_{i+1/2}}\bm{u}\bm{\cdot}\mathrm{d}\bm{S},$ (3)
where $\mathrm{d}\bm{S}$ is the measure denoting integration over the facet. When the mesh is one-dimensional, $V_{i}=\Delta x_{i}$ and $S_{i+1/2}=1$.
@@ 2.1.1 Subgrid Reconstruction
The subgrid reconstruction determines the accuracy of the scheme, both in terms of formal order-of-accuracy and in terms of monotonicity or positivity. A constant reconstruction results in a first-order scheme, whereas a linear reconstruction, such as that of Fromm [1968] and Van Leer [1974], is second-order. The formally third-order quadratic reconstruction, such as used in the NIRVANA scheme of Leonard et al. [1995] or the Piecewise Parabolic Method of Colella and Woodward [1984], PPM (see Appendix A.1 for the discretization used in this paper), is of the form
@@ 2.1.2 Departure Point Calculation
For non-uniform grids, $c_{i+1/2}$ must be computed so that the integer and fractional parts have the same meaning as for the uniform grid. The integer part is thus computed as
@@
where the volume factor in the denominator is the volume of the cell containing the departure point.
@@ 2.1.3 Flux Integral
Given the departure point $x^{dep}_{i+1/2}$, the flux is computed as the integral of the mass between the flux point and the departure point as
$F_{i+1/2}=\frac{1}{S_{i+1/2}\Delta{t}}\int^{V(x=x_{i+1/2})}_{V(x=x^{dep}_{i+1/2})}q\hskip 2.13394pt\mathrm{d}{V}.$ (12)
@@
$F^{int}_{i+1/2}=\begin{dcases}\frac{1}{S_{i+1/2}\Delta{t}}\sum_{j=1}^{c_{i+1/2}^{int}}q_{i+1-j}V_{i+1-j},&u_{i+1/2}\geq 0,\\ \frac{1}{S_{i+1/2}\Delta{t}}\sum_{j=1}^{|c_{i+1/2}^{int}|}q_{i+j}V_{i+j},&u_{i+1/2}<0.\end{dcases}$ (13)
The fractional part uses the subgrid reconstruction from the chosen scheme:
@@
$F_{i+1/2}^{frac}=\widehat{u}_{i+1/2}^{frac}\mathcal{R}(q,c_{i+1/2}^{frac}),$ (15)
where $\mathcal{R}(q,c^{frac})$ can be interpreted as a reconstruction of $q$ at the facet. Importantly, if $q$ is some constant $K$ then $\mathcal{R}(q,c_{i+1/2}^{frac})=K$ irrespective of $c_{i+1/2}^{frac}$. Appendix A discusses the computation of $\mathcal{R}$ for PPM.
@@ 2.1.4 One-Dimensional Scheme for Density
The one-dimensional scheme for density transport (in the $x$-direction) can then be summarised as a simple one-stage forward-in-time scheme:
$\rho^{n+1}=\rho^{n}-\Delta t\nabla_{x}\cdot\mathcal{F}^{x}(\rho^{n},u),$ (16)
where $\mathcal{F}^{x}(\rho,u)$ is the operation to evaluate the flux $F^{x}$ in the $x$-direction, from the density $\rho$ and velocity $u$. The superscript $n$ denotes the value at the $n$-th time level, while $\nabla_{x}\cdot F^{x}$ is the one-dimensional divergence calculation. At a given point, the evolution of the density is expressed as
$\rho_{i}^{n+1}=\rho^{n}_{i}-\frac{\Delta t}{V_{i}}\left(F^{x}_{i+1/2}S_{i+1/2}-F^{x}_{i-1/2}S_{i-1/2}\right).$ (17)

### v2

@@ 1.1 Desirable Properties for Transport Schemes
Numerical methods for solving the transport equation are a core component of atmospheric models that are used for weather prediction and climate research. These equations typically take two forms:
@@
which are known as the conservative form (1a) for a density field $\rho$ or the advective form (1b) which may be used for instance for a tracer mixing ratio [^1] The mixing ratio of a species is defined as the ratio of the mass density of the species to some reference mass density, which in this work is denoted by $\rho$. In atmospheric applications this could be the density of dry air. $m$. In both forms, $\bm{u}$ is the wind velocity field. Considering the equation for the product $\rho m$ (which describes the density of the tracer species), it can be shown that this obeys:
@@
The relevance of this form of the transport equation is that discretisations of (1a) and (2) provide local (and hence global) mass conservation, in that each cell has a closed mass budget when taking into account the mass fluxes into neighbouring cells. As discussed by Thuburn [2008], these conservation properties are desirable for weather and particularly climate models, so that the discretisation itself does not become a net source or sink of the tracer. Locally-conservative schemes naturally provide a closed budget for the tracer species. Additionally, for the transport of a density, Lin and Rood [1996] and Melvin et al. [2024] argue that any conservative transport scheme should provide increments that are linear in $\bm{\nabla\cdot u}$ when the density is constant, a property that we will describe here as divergence-linearity. The corollary of this is that any density field remains constant under divergence-free flow. [^2] This property is described as “constancy” by Leonard et al. [1996], while it is a sub-property of the “consistency” described by Lin and Rood [1996].
@@
However other properties are also important for tracer transport schemes: it can be vital that the scheme is positivity-preserving so that unphysical negative concentrations are not produced. A stronger possible requirement is that of monotonicity, which prevents the introduction of new minima or maxima in the solution. Note that as discussed by Lauritzen et al. [2011], in contrast to $m$, the product $\rho m$ will not in general remain bounded by its initial extrema in divergent flows, but its evolution should be shape-preserving in the sense that no new unphysical extrema are generated. In this work we focus on the monotonicity requirement of keeping the minima and maxima of mixing ratios bounded by their initial values. It is also important for a conservative tracer transport scheme to be consistent with both the advective form of the equation and the conservative transport of $\rho$ (see [Jöckel et al., 2001] and [Zhang et al., 2008]). Thus the property of consistency is defined as the preservation of a constant mixing ratio, independent of the flow. Finally, many operational numerical weather prediction models use large time steps. It is therefore highly attractive to use a scheme which is stable at large Courant numbers [^3] For instance, the finest resolution global configurations discussed in Walters et al. [2019] correspond to Courant numbers of around 5 in the mid-latitude jets. whilst maintaining computational efficiency and without compromising the other properties listed here. A complete discussion of properties that are desirable for transport schemes can be found in Lauritzen et al. [2011].
@@
In addition to the desirable properties listed above, the transport scheme must work within the parameters of the atmospheric dynamical core, which may result in additional requirements for the transport scheme. Many dynamical cores (e.g. [Chen et al., 2008, Wood et al., 2014, Girard et al., 2014, Melvin et al., 2024]) use hexahedral cells and make use of C-grid staggering in the horizontal and Charney-Phillips staggering [Charney and Phillips, 1953] in the vertical, primarily due to the favourable representation of gravity wave dispersion (see for instance Staniforth and Thuburn [2012] and Thuburn [2022]). With these staggerings, the density $\rho$ of dry air is represented at cell centres, the wind velocity $\bm{u}$ is described at the centres of cell faces and the potential temperature $\theta$ is co-located with the vertical component of the wind velocity (at the centre of the top/bottom faces of cells). It is common for tracers such as moisture mixing ratios to be co-located with the potential temperature (or other entropy-type variable) as this is beneficial for capturing the latent heat exchanges associated with phase changes [Bendall et al., 2023]. This work considers a transport scheme that can be used with these staggerings.
@@ 1.2 Background to Flux-Form Semi-Lagrangian Schemes
Some notable conservative transport methods which meet many of the criteria above include: the CSLAM scheme of Lauritzen et al. [2010] and Harris et al. [2011]; the SLICE scheme of Zerroukat et al. [2002, 2004] and Zerroukat and Allen [2012]; the multi-moment scheme of Tang et al. [2022]; and the substepped finite volume Method-of-Lines method described in Melvin et al. [2024]. However, the focus in this work is another class of schemes that possesses many of the desirable properties: Flux-Form Semi-Lagrangian (FFSL) schemes. These schemes originate from the flux integration methods of Van Leer [1974], Colella and Woodward [1984], Carpenter Jr et al. [1990] and were developed in multiple dimensions by Leonard et al. [1996] and Lin and Rood [1996]. Thanks to their conservation and stability properties, FFSL schemes have been successfully used for transport in geophysical fluid models [Lin, 2004, Putman and Lin, 2007, Zhou et al., 2012, Neale et al., 2013, Gillibrand and Herzfeld, 2016, Harris et al., 2021, Zhang et al., 2023, Mouallem et al., 2023]. FFSL schemes involve computing mass fluxes by integrating fields between flux points and corresponding departure points. Conservation is inherent, as increments are evaluated from the divergence of fluxes, which are shared between cells. Like standard semi-Lagrangian schemes, stability is not conditional on the Courant number which permits large time steps.
Instead, the stability of the scheme depends upon the Lipschitz number, which measures the one-dimensional divergence of the velocity $\bm{u}$ seen by a discretisation with time step $\Delta t$. In other words $\Delta t|\partial u/\partial x|$ in the $x$-direction and similar for other directions. In an Arakawa C-grid staggering, in which density fields are located at cell centres and the velocity is described by its components normal to each facet, the Lipschitz numbers are naturally calculated at cell centres. However Lipschitz numbers can also be calculated at facets, so that in one-dimension $\ell=(c-c^{\dagger})\,\mathrm{sign}(c)$, where $c$ is the Courant number at that facet and $c^{\dagger}$ is the Courant number at the next facet upwind. The result is that when this number is above 1, the trajectories of discrete particles can cross [Smolarkiewicz and Pudykiewicz, 1992].
The multi-dimensional forms of FFSL presented by Leonard et al. [1996] and Lin and Rood [1996] are built from one-dimensional calculations through dimensional-splitting techniques. Following Leonard et al. [1996], the multi-dimensional splitting of Leonard et al. [1996] and Lin and Rood [1996] will be referred to as the COSMIC splitting throughout this work. Advantages of using multiple one-dimensional methods over pure multi-dimensional methods include the ease of applying limiters, the ease of extending to large Courant number flow, and that in general one-dimensional methods are computationally cheaper than their multi-dimensional counterparts.
@@ 2.1 Review of One Dimensional Flux-Form Semi-Lagrangian Schemes
In this section, the one-dimensional FFSL calculation is reviewed, but on a general one-, two- or three-dimensional mesh. At first, calculations of mass fluxes are described for a density $\rho$. For simplicity, the one-dimensional scheme is described in the $x$ direction only, where $x_{i}$ are the coordinates of cell centres, $i$ is the grid index, $V_{i}$ is the volume of the $i$-th cell, and $\rho_{i}=\rho(x_{i})$ is the density in cell $i$. Mass fluxes are computed at cell facets, and denoted by $F_{i+1/2}$. The cell facets $\varGamma_{i+1/2}$ are vertices for one-dimensional meshes, edges for two-dimensional meshes, and faces for three-dimensional meshes, and are indexed by an additional $1/2$ to indicate that these are staggered from cell centres. While $\bm{u}$ is the general wind velocity vector, $u_{i+1/2}$ is its projection onto the normal $\bm{n}$ to the facet $\varGamma_{i+1/2}$. The facet has length or area $S_{i+1/2}$, so that
$u_{i+1/2}S_{i+1/2}=\int_{\varGamma_{i+1/2}}\bm{u\cdot n}\ \mathrm{d}S=\int_{\varGamma_{i+1/2}}\bm{u}\bm{\cdot}\mathrm{d}\bm{S},$ (3)
where $\mathrm{d}\bm{S}$ is the measure denoting integration over the facet. When the mesh is one-dimensional, $V_{i}=\Delta x_{i}$ and $S_{i+1/2}=1$. For a multi-dimensional mesh, $\bm{u}$ can naturally be expanded through either primal basis vectors (which point tangential to mesh lines) or dual basis vectors (which are normal to mesh lines). These are not equivalent if the mesh is not orthogonal, in which case the quantity $u_{i+1/2}$ corresponds to the contravariant component of $\bm{u}$, which multiplies the primal basis vector on $\varGamma_{i+1/2}$.
@@ 2.1.1 Subgrid Reconstruction
The subgrid reconstruction determines the accuracy of the scheme, both in terms of formal order-of-accuracy and in terms of monotonicity or positivity. A constant reconstruction results in a first-order scheme, whereas a linear reconstruction, such as that of Fromm [1968] and Van Leer [1974], is second-order. The formally third-order quadratic reconstruction, such as used in the NIRVANA scheme of Leonard et al. [1995] or the Piecewise Parabolic Method of Colella and Woodward [1984], PPM (see Appendix A for the discretization used in this paper), is of the form
@@ 2.1.2 Departure Point Calculation
For a general non-uniform grid, $c_{i+1/2}$ must be computed so that the integer and fractional parts have the same meaning as for the uniform grid. The integer part is thus computed as
@@
where the volume factor in the denominator is the volume of the cell containing the departure point. Note that in some of the following sections, the volume factors appearing in (9), (10) and (11) will be substituted for other fields, such as a mass field.
@@ 2.1.3 Flux Integral
Given the departure point $x^{dep}_{i+1/2}$, the flux of $\rho$ is computed as the integral of the mass between the flux point and the departure point as
$F_{i+1/2}=\frac{1}{S_{i+1/2}\Delta{t}}\int^{V(x=x_{i+1/2})}_{V(x=x^{dep}_{i+1/2})}\rho\hskip 2.13394pt\mathrm{d}{V}.$ (12)
@@
$F^{int}_{i+1/2}=\begin{dcases}\frac{1}{S_{i+1/2}\Delta{t}}\sum_{j=1}^{c_{i+1/2}^{int}}\rho_{i+1-j}V_{i+1-j},&u_{i+1/2}\geq 0,\\ \frac{1}{S_{i+1/2}\Delta{t}}\sum_{j=1}^{|c_{i+1/2}^{int}|}\rho_{i+j}V_{i+j},&u_{i+1/2}<0.\end{dcases}$ (13)
The fractional part uses the subgrid reconstruction from the chosen scheme. Although any reconstruction can be used, in this work we focus on the quadratic PPM reconstruction of $\mathcal{Q}(\xi)$ as in (4). The fractional flux integrates the subgrid reconstruction from the facet to the departure point:
@@
$F_{i+1/2}^{frac}=\widehat{u}_{i+1/2}^{frac}\mathcal{R}(\rho,c_{i+1/2}^{frac}),$ (15)
where $\mathcal{R}(\rho,c^{frac})$ can be interpreted as a reconstruction of $\rho$ at the facet. Importantly, if $\rho$ is some constant $K$ then $\mathcal{R}(\rho,c_{i+1/2}^{frac})=K$ irrespective of $c_{i+1/2}^{frac}$. Appendix A discusses the computation of $\mathcal{R}$ for PPM.
If the density $\rho$ is a constant $K$ across the domain, then the mass flux is simply $F_{i+1/2}=Ku_{i+1/2}$, which can be seen by considering the evaluation of the integer and fractional parts of the flux. The integer part of the flux simplifies to
$F^{int}_{i+1/2}=\begin{dcases}\frac{K}{S_{i+1/2}\Delta{t}}\sum_{j=1}^{c_{i+1/2}^{int}}V_{i+1-j},&u_{i+1/2}\geq 0,\\ \frac{K}{S_{i+1/2}\Delta{t}}\sum_{j=1}^{|c_{i+1/2}^{int}|}V_{i+j},&u_{i+1/2}<0,\end{dcases}$ (16)
which can be compared with (14) to show that
$F^{int}_{i+1/2}=K\left(u_{i+1/2}-\widehat{u}^{frac}_{i+1/2}\right).$ (17)
Using (15) with a constant $\rho$, the fractional flux is
$F^{frac}_{i+1/2}=K\widehat{u}^{frac}_{i+1/2},$ (18)
and therefore
$F_{i+1/2}=F^{int}_{i+1/2}+F_{i+1/2}^{frac}=K\left(u_{i+1/2}-\widehat{u}^{frac}_{i+1/2}\right)+K\widehat{u}^{frac}_{i+1/2}=Ku_{i+1/2}.$ (19)
@@ 2.1.4 One-Dimensional Scheme for Density
The procedure described in Sections 2.1.2 and 2.1.3 is used to compute departure displacements and then mass fluxes at each cell facet. The density is then evolved in time through a simple forward-in-time scheme, so that at a given point:
$\rho_{i}^{n+1}=\rho^{n}_{i}-\frac{\Delta t}{V_{i}}\left(F^{x}_{i+1/2}S_{i+1/2}-F^{x}_{i-1/2}S_{i-1/2}\right),$ (20)
where the superscript $n$ denotes the density at the $n$-th time level.
For the remainder of this work, we use the operator
$\mathcal{F}^{x}(\rho,u^{x},V)$ (21)
to describe the process of computing departure displacements and evaluating mass fluxes (in the $x$-direction) for the whole domain from the density $\rho$, velocity component $u^{x}$ and cell volume $V$, following the procedure in Sections 2.1.2 and 2.1.3. The final argument represents each time that the cell volume appears in the computation of departure displacements and mass fluxes. With this notation, the one-dimensional scheme for density transport (in the $x$-direction) can be summarised as:
$\rho^{n+1}=\rho^{n}-\Delta t\nabla_{x}\cdot\mathcal{F}^{x}(\rho^{n},u^{x},V),$ (22)
where $\nabla_{x}\cdot$ indicates the one-dimensional discrete divergence calculation. When $\rho^{n}$ is a constant $K$, the mass flux reduces to $Ku^{x}$ and so the value of the density at the next time level is $K(1-\Delta t\nabla_{x}\cdot u^{x})$. This satisfies the property of divergence-linearity.